Find an equation for the line with the given properties:
Passing thru the point (-4,-1), parallel to the line x - 4y = 2
I wrote x - 4y = 2 in slope intercept form, as y = -1/2x + 1/4
is this correct?

Answer by mmm4444bot(95)

Hello There:
Your equation in slope-intercept form is not correct.
x - 4*y = 2
Subtract x from both sides.
-4*y = -x + 2
Divide both sides by -4.
y = (1/4)*x - (1/2)
You can now use the point-slope equation with the coordinates of (-4, -1) and the slope of 1/4 to find the equation this problem asks for.
~ Mark